Circular No. 7257

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GM SAGITTARII AND SAX J1819.3–2525 = XTE J1819–254 M. L. McCollough and M. H. Finger, Universities Space Research As-

sociation and Marshall Space Flight Center; and P. M. Woods, University of Alabama, Huntsville, report for the BATSE team: “GM Sgr appears as a strong and highly variable source in the BATSE energy range. The recent x-ray event (IAUC 7253) was first detected in the BATSE data on Sept. 14.88 UT with a 20–100-keV flux of ∼ 5 Crab (9.0 × 10^{−8} erg cm^{−2} s^{−1} with a photon index of −3.5). The occultation measurement prior to this showed no excess flux above background. From Sept. 15.36 to 15.68, the source was detected at fluxes reaching ∼ 8 Crab (20–100 keV). Power spectra made from 20–50-keV rates during this interval were all consistent with a featureless 1/f continuum from 5 to 500 mHz. The fractional rms amplitudes for the range 10–500 mHz varied from 17 to 38 percent. For the remainder of the day, the source was undetected except for several short flares of ∼ 10–500 s. The peak flux values of these events reach 1.2 × 10^{−7} erg cm^{−2} s^{−1} (20–100 keV), with the source being clearly detected above 100 keV.”

C. B. Markwardt, University of Maryland and Goddard Space Flight Center (GSFC); J. H. Swank, GSFC; and E. H. Morgan, Massachusetts Institute of Technology, report that the RXTE PCA observed the transient source XTE J1819–254 starting on Sept. 15.89 UT for ∼ 3000 s. The intensity of the source during the observation ranged from (1.8 to 860) × 10^{−10} erg s^{−1} cm^{−2} in the band 2–60 keV. After a flaring period, characterized by rapid shots and lasting ∼ 1000 s, it entered a quiescent phase at the minimum flux level. The x-ray spectrum during the flare was consistent with a power law having a photon index of 0.9, a rolloff above 20 keV, a soft excess, and neutral hydrogen absorption < 2 × 10^{21} cm^{−2}. During the quiescent phase, the spectrum steepened to a photon index of 3. An iron emission line is present in both spectra, with a best-fit centroid of 6.5 keV and equivalent width of ∼ 500 eV. No significant pulsations or quasiperiodic variability were detected, although red noise was present below 30 Hz. Subsequent PCA observations on Sept. 16.06, 16.12, and 16.18 also showed quiescent behavior at similar intensities. We have also undertaken regular monitoring observations of the galactic-center region. These show that the source has varied irregularly in the band 2–60 keV between (0.2 and 12) × 10^{−10} erg s^{−1} cm^{−2} over the past 7 months, the most recent measurements being (in the same units) Sept. 9.30, 2.1; 12.15, 3.1; 15.10, 4.1. Additional public RXTE PCA observations are planned.

1999 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green